Thursday 10/18/12

9:40 A – 7:00P, 8:20P – 10:45P

  • Flip collection cage, 9:45A.  Keep plate to make new bottles
  • Pick up more 10x TAE and 20x SSC from stock room
  • Picked up 6 new pints of 100% EtOH.
  • Found issue with K27:H3 ratio.  Fixed = much cleaner data!
DNA FISH probe making
  • Spin down BAC7 DNA probes (slightly colored pellets. sample 2 (‘false start’) has more blue color than sample 1.
  • resuspend in D-FISH hybe buffer
  •  run 5 uL on a gel
  • No probe — sample not substantially digested.
  • Ordered new DNA pol I and DNAse I components from Promega following JW protocol recommendations.
  • Test synthesis with regular DNTPs and a range of DNAse concentrations (2,4,6,10,14 uL in 50 uL rxn).  All with 1 ug of BAC 7 DNA.
  • Running test gel:  Bands from original DNA all visible.  Possibly slightly fainter at higher DNAse concentration but not dramatically so.
Embryo labeling
  • rehydrating embryos
Fly work
  • Flipped MTD x Pc, MTD x Psc, and MTD x Scm into new bottles.  Embryos/larvae growing well.
  • Flipped sim[D] into 2 new bottles.   Should be ready for collection cage next generation
  • Flipped yw Act5-Gal4 into new bottle.  
  • Moved sna[18], Sp/CyoZ and Pr/HbZ into bottles.  need to start building double-balanced sna[18] again (previous one died and Mounia hasn’t been able to do it).  

embryo fixation
  • Thawing DFISH mix
  • # PM Fix embryos
  • # Block 1 hr
  • # stain with rb anti-Pc O/N at 4C

Results to date on K27me3 dynamics

  • found bug in analysis of K27me3 to H3 rato, now fixed (?).
  • Fix image 5 nuclear mask (others look okay);
This entry was posted in Summaries and tagged , . Bookmark the permalink.